OpenStudy (anonymous):

x intercept of 3(2-x^1/2)/x

OpenStudy (anonymous):

Is your question right ??

OpenStudy (anonymous):

Is it :

OpenStudy (anonymous):

yes do u know wat 1/2 power is

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

It is simply square root bracket..

OpenStudy (anonymous):

did i win the gold :)

OpenStudy (anonymous):

yes it is 4 how u get it

OpenStudy (anonymous):

y = 3(2 - root(x))/x Put y = 0 So: 3(2 - root(x)) = 0 or: 2 - root(x) = 0 root(x) = 2 Square both the sides: x = 2^2 = 4..

OpenStudy (anonymous):

how u lose 3 thats where i get lost

OpenStudy (anonymous):

Divide by 3 both the sides..

OpenStudy (anonymous):

ok thanks how do i gove medals

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

thank you :)